# Scaling Depth Can Enable New Goal-Reaching Capabilities

### v2

@@ Abstract
Scaling up self-supervised learning has driven breakthroughs in language and vision, yet comparable progress has remained elusive in reinforcement learning (RL). In this paper, we study building blocks for self-supervised RL that unlock substantial improvements in scalability, with network depth serving as a critical factor. Whereas most RL papers in recent years have relied on shallow architectures (around 2 – 5 layers), we demonstrate that increasing the depth up to 1024 layers can significantly boost performance. Our experiments are conducted in an unsupervised goal-conditioned setting, where no demonstrations or rewards are provided, so an agent must explore (from scratch) and learn how to maximize the likelihood of reaching commanded goals. Evaluated on simulated locomotion and manipulation tasks, our approach increases performance by 2×2$\times$2 ×50×50$\times$50 ×. Increasing the model depth not only increases success rates but also qualitatively changes the behaviors learned.
reinforcement learning, scaling, self-supervised learning, contrastive learning
00footnotetext: Website + code: https://wang-kevin3290.github.io/scaling-crl/
Figure 1: Scaling network depth yields performance gains across a suite of locomotion, navigation, and manipulation tasks, ranging from doubling performance to 50× improvements on Humanoid-based tasks. Notably, rather than scaling smoothly, performance often jumps at specific critical depths (e.g., 8 layers on Ant Big Maze, 64 on Humanoid U-Maze), which correspond to the emergence of qualitatively distinct learned policies (see Section 4).
@@ 1 Introduction
In this paper, we will study building blocks for scaling reinforcement learning. Our first step is to rethink the conventional wisdom above: “reinforcement learning” and “self-supervised learning” are not diametric learning rules, but rather can be married together into self-supervised RL systems that explore and learn policies without reference to a reward function or demonstrations (Eysenbach et al.,, 2021, 2022; Lee et al.,, 2022) . In this work, we use one of the simplest self-supervised RL algorithms, contrastive RL (CRL) (Eysenbach et al.,, 2022). The second step is to recognize the importance of increasing available data. We will do this by building on recent GPU-accelerated RL frameworks (Makoviychuk et al.,, 2021; Rutherford et al.,, 2023; Rudin et al.,, 2022; Bortkiewicz et al.,, 2024). The third step is to increase network depth, using networks that are up to 100×100$\times$100 × deeper than those typical in prior work. Stabilizing the training of such networks will require incorporating architectural techniques from prior work, including residual connections (He et al.,, 2015), layer normalization (Ba et al.,, 2016), and Swish activation (Ramachandran et al.,, 2018). Our experiments will also study the relative importance of batch size and network width.
@@
  • Empirical Scalability: We observe a significant performance increase, more than 20×20$\times$20 × in half of the environments, corresponding to qualitatively distinct policies that emerge with scale.
@@
  • Empirical Analysis: We show that deeper networks exhibit improved stitching capabilities, learn more accurate value functions, and unlock the benefits of larger batch sizes.
@@ 2 Related Work
Recent approaches suggest several promising directions, including new architectural paradigms (Obando-Ceron et al.,, 2024), distributed training approaches (Ota et al.,, 2021; Espeholt et al.,, 2018), distributional RL (Kumar et al.,, 2023), and distillation (Team et al.,, 2023). Compared to these approaches, our method makes a simple extension to an existing self-supervised RL algorithm. The most recent works in this vein include Lee et al., (2024) and Nauman et al., 2024b , which leverage residual connections to facilitate the training of wider networks. These efforts primarily focus on network width, noting limited gains from additional depth, thus both works use architectures with only four MLP layers. In our experiments, we find that scaling width indeed improves performance (Section 4.2); however, our approach also enables scaling along depth, proving to be more powerful than width alone.
@@ 3 Preliminaries
Our experiments will use a contrastive RL algorithm (Eysenbach et al.,, 2022) to solve goal-conditioned problems. Contrastive RL is an actor-critic method; we will use $f_{\phi,\psi}(s,a,g)$ to denote the critic and $\pi_{\theta}(a\mid s,g)$ to denote the policy. The critic is parametrized with two neural networks that return state, action pair embedding $\phi(s,a)$ and goal embedding $\psi(g)$. The critic’s output is defined as the $l^{2}$-norm between these embeddings: $f_{\phi,\psi}(s,a,g)=\|\phi(s,a)-\psi(g)\|_{2}$. The critic is trained with the InfoNCE objective (Sohn,, 2016) as in previous works (Eysenbach et al.,, 2022, 2021; Zheng et al.,, 2023, 2024; Myers et al.,, 2024; Bortkiewicz et al.,, 2024). Training is conducted on batches $\mathcal{B}$, where ${\color[rgb]{0,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.55,0}s_{i},a_{i},g_{i}}$ represent the state, action, and goal (future state) sampled from the same trajectory, while ${\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g_{j}}$ represents a goal sampled from a different, random trajectory. The objective function is defined as:
@@
Figure 2: Architecture. Our approach integrates residual connections into both the actor and critic networks of the Contrastive RL (CRL) algorithm. The depth of this residual architecture is defined as the total number of Dense layers across the residual blocks, which with our residual block size of 4, equates to $4N$.
@@ 4 Experiments
In this section, we detail the experiments conducted to evaluate the scalability and effectiveness of our approach across a diverse set of tasks and configurations. We describe the experimental setup before presenting the key empirical results, analyses, and ablations.  [^1] Project website with code: https://wang-kevin3290.github.io/scaling-crl/
@@ 4.1 Experimental Setup
All RL experiments use the JaxGCRL codebase (Bortkiewicz et al.,, 2024), which facilitates fast online GCRL experiments based on Brax (Freeman et al.,, 2021) and MJX (Todorov et al.,, 2012) environments. The specific environments used are a range of locomotion, navigation, and robotic manipulation tasks, as shown in Figure 3.
Figure 3: The scaling results of this paper are demonstrated on the JaxGCRL benchmark, showing that they replicate across a diverse range of locomotion, navigation, and manipulation tasks. These tasks are set in the goal-conditioned setting where there are no auxiliary rewards or demonstrations. Figure taken from (Bortkiewicz et al.,, 2024).
While no rewards are used during training, for evaluation we will measure the number of time steps (out of 1000) that the agent is near the goal. When reporting an algorithm’s performance as a single number, we will compute the average score over the last five epochs of training.
@@
We employ residual connections from the ResNet architecture (He et al.,, 2015), with each residual block consisting of four repeated units of a Dense layer, a Layer Normalization (Ba et al.,, 2016) layer, and Swish activation (Ramachandran et al.,, 2018). We apply the residual connections immediately following the final activation of the residual block, as shown in Figure 2. In this paper, we define the depth of the network as the total number of Dense layers across all residual blocks in the architecture. In all experiments, the depth refers to the configuration of the actor network and both critic encoder networks, which are scaled jointly, except for the ablation experiment in Section 4.9.
@@ 4.2 Scaling Depth in Contrastive RL
Table 1: Increasing network depth (depth $D=4\rightarrow 64$) increased performance (Figure 1). Scaling depth exhibits the greatest benefits on tasks with the largest observation dimension (Dim).
Task Dim D = 4 D = 64 Imprv.
Arm Binpick Hard 17 $38$$\mathbf{\pm}4$ $219$$\mathbf{\pm}15$ 5.7×5.7$\times$5.7 ×
Arm Push Easy $308$$\mathbf{\pm}33$ $762$$\mathbf{\pm}30$ 2.5×2.5$\times$2.5 ×
Arm Push Hard $171$$\mathbf{\pm}11$ $410$$\mathbf{\pm}13$ 2.4×2.4$\times$2.4 ×
Ant U4-Maze 29 $11.4$$\mathbf{\pm}4.1$ $286$$\mathbf{\pm}36$ 25×25$\times$25 ×
Ant U5-Maze $0.97$$\mathbf{\pm}0.7$ $61$$\mathbf{\pm}18$ 63×63$\times$63 ×
Ant Big Maze $61$$\mathbf{\pm}20$ $441$$\mathbf{\pm}25$ 7.3×7.3$\times$7.3 ×
Ant Hardest Maze $215$$\mathbf{\pm}8$ $387$$\mathbf{\pm}21$ 1.8×1.8$\times$1.8 ×
Humanoid 268 $12.6$$\mathbf{\pm}1.3$ $649$$\mathbf{\pm}19$ 52×52$\times$52 ×
Humanoid U-Maze $3.2$$\mathbf{\pm}1.2$ $159$$\mathbf{\pm}33$ 50×50$\times$50 ×
Humanoid Big Maze $0.06$$\mathbf{\pm}0.04$ $59$$\mathbf{\pm}21$ 1051×1051$\times$1051 ×
We start by studying how increasing network depth can increase performance. Both the JaxGCRL benchmark and relevant prior work (Lee et al.,, 2024; Nauman et al., 2024b, ; Zheng et al.,, 2024) use MLPs with a depth of 4, and as such we adopt it as our baseline. In contrast, we will study networks of depth 8, 16, 32, and 64. The results in Figure 1 demonstrate that deeper networks achieve significant performance improvements across a diverse range of locomotion, navigation, and manipulation tasks. Compared to the 4-layer models typical in prior work, deeper networks achieve 25×2-5$\times$2 - 5 × gains in robotic manipulation tasks, over 20×20$\times$20 × gains in long-horizon maze tasks such as Ant U4-Maze and Ant U5-Maze, and over 50×50$\times$50 × gains in humanoid-based tasks. The full table of performance increases up to depth 64 is provided in Table 1.
Figure 4: Increasing depth results in new capabilities:
Row 1
: A humanoid agent trained with network depth of 4 collapses and throws itself towards the goal, as opposed to in Row 2, where the depth 16 agent gains the ability to walk upright. Row 3: At depth 64, the humanoid agent in U-Maze struggles to reach the goal and falls. Row 4: A novel policy emerges at depth 256, as the agent exhibits an acrobatic strategy of compressing its body to vault over the maze wall.
@@ 4.3 Emergent Policies Through Depth
A closer examination of the results from the performance curves in Figure 1 reveals a notable pattern: instead of a gradual improvement in performance as depth increases, there are pronounced jumps that occur once a critical depth threshold is reached (also shown in Figure 9). The critical depths vary by environment, ranging from 8 layers (e.g. Ant Big Maze) to 64 layers in the Humanoid U-Maze task, with further jumps occurring even at depths of 1024 layers (see the Testing Limits section, Section 4.6).
@@
Prompted by this observation, we visualized the learned policies at various depths and found qualitatively distinct skills and behaviors exhibited. This is particularly pronounced in the humanoid-based tasks, as illustrated in Figure 4. Networks with a depth of 4 exhibit rudimentary policies where the agent either falls or throws itself toward the target. Only at a critical depth of 16 does the agent develop the ability to walk upright into the goal. In the Humanoid U-Maze environment, networks of depth 64 struggle to navigate around the intermediary wall, collapsing on the ground. Remarkably at a depth of 256, the agent learns unique behaviors on Humanoid U-Maze. These behaviors include folding forward into a leveraged position to propel itself over walls and shifting into a seated posture over the intermediary obstacle to worm its way toward the goal (one of these policies is illustrated in the fourth row of Figure 4). To the best of our knowledge, this is the first goal-conditioned approach to document such behaviors on the humanoid environment.
@@
4.4 Depth Enhances Contrastive Representations
The long-horizon setting has been a long-standing challenge in RL particularly in unsupervised goal-conditioned settings where there is no auxiliary reward feedback (Gupta et al.,, 2019). The family of U-Maze environments requires a global understanding of the maze layout for effective navigation. We consider a variant of the Ant UMaze environment, the U4-maze, in which the agent must initially move in the direction opposite the goal to loop around and ultimately reach it. As shown in Figure 5, we observe a qualitative difference in the behavior of the shallow network (depth 4) compared to the deep network (depth 64). The visualized Q-values computed from the critic encoder representations reveal that the depth 4 network seemingly relies on Euclidean distance to the goal as a proxy for the Q value, even when a wall obstructs the direct path. In contrast, the depth 64 critic network learns richer representations, enabling it to effectively capture the topology of the maze as visualized by the trail of high Q values along the inner edge. These findings suggest that scaling network depth may be an important ingredient for reasoning over longer horizons in sparse reward settings.
4.5 Deeper Networks Enable Partial Experience Stitching
@@
Another key challenge in reinforcement learning is learning policies that can generalize to tasks unseen during training. To evaluate this setting, we designed a modified version of the Ant U-Maze environment. As shown in Figure 6 (top right), the original JaxGCRL benchmark assesses the agent’s performance on the three farthest goal positions located on the opposite side of the wall. However, instead of training on all possible subgoals (a superset of the evaluation state-goal pairs), we modified the setup to train on start-goal pairs that are at most 3 units apart, ensuring that none of the evaluation pairs ever appear in the training set. Figure 6 demonstrates that depth 4 networks show limited generalization, solving only the easiest goal (4 units away from the start). Depth 16 networks achieve moderate success, while depth 64 networks excel, sometimes solving the most challenging goal position. These results suggest that the increasing network depth results in some degree of stitching, combining $\leq$3-unit pairs to navigate the 6-unit span of the U-Maze. These results demonstrate that scaling depth can improve generalization.

### v3

# Kevin Wang

@@ Abstract
Scaling up self-supervised learning has driven breakthroughs in language and vision, yet comparable progress has remained elusive in reinforcement learning (RL). In this paper, we study building blocks for self-supervised RL that unlock substantial improvements in scalability, with network depth serving as a critical factor. Whereas most RL papers in recent years have relied on shallow architectures (around 2 – 5 layers), we demonstrate that increasing the depth up to 1024 layers can significantly boost performance. Our experiments are conducted in an unsupervised goal-conditioned setting, where no demonstrations or rewards are provided, so an agent must explore (from scratch) and learn how to maximize the likelihood of reaching commanded goals. Evaluated on simulated locomotion and manipulation tasks, our approach increases performance on the self-supervised contrastive RL algorithm by $2\times$$50\times$, outperforming other goal-conditioned baselines. Increasing the model depth not only increases success rates but also qualitatively changes the behaviors learned. The project webpage and code can be found here: https://wang-kevin3290.github.io/scaling-crl/.
Figure 1: Scaling network depth yields performance gains across a suite of locomotion, navigation, and manipulation tasks, ranging from doubling performance to 50× improvements on Humanoid-based tasks. Notably, rather than scaling smoothly, performance often jumps at specific critical depths (e.g., 8 layers on Ant Big Maze, 64 on Humanoid U-Maze), which correspond to the emergence of qualitatively distinct policies (see Section 4).
@@ 1 Introduction
In this paper, we will study building blocks for scaling reinforcement learning. Our first step is to rethink the conventional wisdom above: “reinforcement learning” and “self-supervised learning” are not diametric learning rules, but rather can be married together into self-supervised RL systems that explore and learn policies without reference to a reward function or demonstrations (Eysenbach et al., , 2021, 2022; Lee et al., , 2022). In this work, we use one of the simplest self-supervised RL algorithms, contrastive RL (CRL) (Eysenbach et al., , 2022). The second step is to recognize the importance of increasing available data. We will do this by building on recent GPU-accelerated RL frameworks (Makoviychuk et al., , 2021; Rutherford et al., , 2023; Rudin et al., , 2022; Bortkiewicz et al., , 2024). The third step is to increase network depth, using networks that are up to $100\times$ deeper than those typical in prior work. Stabilizing the training of such networks will require incorporating architectural techniques from prior work, including residual connections (He et al., , 2015), layer normalization (Ba et al., , 2016), and Swish activation (Ramachandran et al., , 2018). Our experiments will also study the relative importance of batch size and network width.
@@
  • Empirical Scalability: We observe a significant performance increase, more than $20\times$ in half of the environments and out-performing other standard goal-conditioned baselines. These performances gains correspond to qualitatively distinct policies that emerge with scale.
@@
  • Empirical Analysis: We conduct an extensive analysis of the key components in our scaling approach, uncovering critical factors and offering new insights.
@@ 2 Related Work
Recent approaches suggest several promising directions, including new architectural paradigms (Obando-Ceron et al., , 2024), distributed training approaches (Ota et al., , 2021; Espeholt et al., , 2018), distributional RL (Kumar et al., , 2023), and distillation (Team et al., , 2023). Compared to these approaches, our method makes a simple extension to an existing self-supervised RL algorithm. The most recent works in this vein include Lee et al., (2024) and Nauman et al., 2024b , which leverage residual connections to facilitate the training of wider networks. These efforts primarily focus on network width, noting limited gains from additional depth, thus both works use architectures with only four MLP layers. In our method, we find that scaling width indeed improves performance (Section˜4.4); however, our approach also enables scaling along depth, proving to be more powerful than width alone.
@@ 3 Preliminaries
Our experiments will use the contrastive RL algorithm (Eysenbach et al., , 2022) to solve goal-conditioned problems. Contrastive RL is an actor-critic method; we will use $f_{\phi,\psi}(s,a,g)$ to denote the critic and $\pi_{\theta}(a\mid s,g)$ to denote the policy. The critic is parametrized with two neural networks that return state, action pair embedding $\phi(s,a)$ and goal embedding $\psi(g)$. The critic’s output is defined as the $l^{2}$-norm between these embeddings: $f_{\phi,\psi}(s,a,g)=\|\phi(s,a)-\psi(g)\|_{2}$. The critic is trained with the InfoNCE objective (Sohn, , 2016) as in previous works (Eysenbach et al., , 2022, 2021; Zheng et al., , 2023, 2024; Myers et al., , 2024; Bortkiewicz et al., , 2024). Training is conducted on batches $\mathcal{B}$, where $\color[rgb]{0,0.55,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.55,0}s_{i},a_{i},g_{i}$} represent the state, action, and goal (future state) sampled from the same trajectory, while $\color[rgb]{0.75,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.75,0,0}g_{j}$} represents a goal sampled from a different, random trajectory. The objective function is defined as:
@@
Figure 2: Architecture. Our approach integrates residual connections into both the actor and critic networks of the Contrastive RL algorithm. The depth of this residual architecture is defined as the total number of Dense layers across the residual blocks, which, with our residual block size of 4, equates to $4N$.
@@ 4.1 Experimental Setup
All RL experiments use the JaxGCRL codebase (Bortkiewicz et al., , 2024), which facilitates fast online GCRL experiments based on Brax (Freeman et al., , 2021) and MJX (Todorov et al., , 2012) environments. The specific environments used are a range of locomotion, navigation, and robotic manipulation tasks, for details see Appendix˜B. We use a sparse reward setting, with $r=1$ only when the agent is in the goal proximity. For evaluation, we measure the number of time steps (out of 1000) that the agent is near the goal. When reporting an algorithm’s performance as a single number, we compute the average score over the last five epochs of training.
@@
We employ residual connections from the ResNet architecture (He et al., , 2015), with each residual block consisting of four repeated units of a Dense layer, a Layer Normalization (Ba et al., , 2016) layer, and Swish activation (Ramachandran et al., , 2018). We apply the residual connections immediately following the final activation of the residual block, as shown in Figure˜2. In this paper, we define the depth of the network as the total number of Dense layers across all residual blocks in the architecture. In all experiments, the depth refers to the configuration of the actor network and both critic encoder networks, which are scaled jointly, except for the ablation experiment in Section˜4.4.
@@ 4.2 Scaling Depth in Contrastive RL
We start by studying how increasing network depth can increase performance. Both the JaxGCRL benchmark and relevant prior work (Lee et al., , 2024; Nauman et al., 2024b, ; Zheng et al., , 2024) use MLPs with a depth of 4, and as such we adopt it as our baseline. In contrast, we will study networks of depth 8, 16, 32, and 64. The results in Figure˜1 demonstrate that deeper networks achieve significant performance improvements across a diverse range of locomotion, navigation, and manipulation tasks. Compared to the 4-layer models typical in prior work, deeper networks achieve $2-5\times$ gains in robotic manipulation tasks, over $20\times$ gains in long-horizon maze tasks such as Ant U4-Maze and Ant U5-Maze, and over $50\times$ gains in humanoid-based tasks. The full table of performance increases up to depth 64 is provided in Table˜1.
In Figure˜12, we present results the same 10 environments, but compared against SAC, SAC+HER, TD3+HER, GCBC, and GCSL. Scaling CRL leads to substantial performance improvements, outperforming all other baselines in 8 out of 10 tasks. The only exception is SAC on the Humanoid Maze environments, where it exhibits greater sample efficiency early on; however, scaled CRL eventually reaches comparable performance. These results highlight that scaling the depth of the CRL algorithm enables state-of-the-art performance in goal-conditioned reinforcement learning.
@@ 4.3 Emergent Policies Through Depth
Figure 3: Increasing depth results in new capabilities: Row 1: A depth-4 agent collapses and throws itself toward the goal. Row 2: A depth-16 agent walks upright. Row 3: A depth-64 agent struggles and falls. Row 4: A depth-256 agent vaults the wall acrobatically.
@@
A closer examination of the results from the performance curves in Figure˜1 reveals a notable pattern: instead of a gradual improvement in performance as depth increases, there are pronounced jumps that occur once a critical depth threshold is reached (also shown in Figure˜6). The critical depths vary by environment, ranging from 8 layers (e.g. Ant Big Maze) to 64 layers in the Humanoid U-Maze task, with further jumps occurring even at depths of 1024 layers (see the Testing Limits section, Section˜4.4).
@@
Prompted by this observation, we visualized the learned policies at various depths and found qualitatively distinct skills and behaviors exhibited. This is particularly pronounced in the humanoid-based tasks, as illustrated in Figure˜3. Networks with a depth of 4 exhibit rudimentary policies where the agent either falls or throws itself toward the target. Only at a critical depth of 16 does the agent develop the ability to walk upright into the goal. In the Humanoid U-Maze environment, networks of depth 64 struggle to navigate around the intermediary wall, collapsing on the ground. Remarkably at a depth of 256, the agent learns unique behaviors on Humanoid U-Maze. These behaviors include folding forward into a leveraged position to propel itself over walls and shifting into a seated posture over the intermediary obstacle to worm its way toward the goal (one of these policies is illustrated in the fourth row of Figure˜3). To the best of our knowledge, this is the first goal-conditioned approach to document such behaviors on the humanoid environment.
4.4 What Matters for CRL Scaling
Width vs. Depth
Figure 4: Scaling network width vs. depth. Here, we reflect findings from previous works (Lee et al., , 2024; Nauman et al., 2024b, ) which suggest that increasing network width can enhance performance. However, in contrast to prior work, our method is able to scale depth, yielding more impactful performance gains. For instance, in the Humanoid environment, raising the width to 2048 (depth=4) fails to match the performance achieved by simply doubling the depth to 8 (width=256). The comparative advantage of scaling depth is more pronounced as the observational dimensionality increases.
Past literature has shown that scaling network width can be effective  (Lee et al., , 2024; Nauman et al., 2024b, ). In Figure˜4, we find that scaling width is also helpful in our experiments: wider networks consistently outperform narrower networks (depth held constant at 4). However, depth seems to be a more effective axis for scaling: simply doubling the depth to 8 (width held constant at 256) outperforms the widest networks in all three environments. The advantage of depth scaling is most pronounced in the Humanoid environment (observation dimension 268), followed by Ant Big Maze (dimension 29) and Arm Push Easy (dimension 17), suggesting that the comparative benefit may increase with higher observation dimensionality.
Note additionally that the parameter count scales linearly with width but quadratically with depth. For comparison, a network with 4 MLP layers and 2048 hidden units has roughly 35M parameters, while one with a depth of 32 and 256 hidden units has only around 2M. Therefore, when operating under a fixed FLOP compute budget or specific memory constraints, depth scaling may be a more computationally efficient approach to improving network performance.
Scaling the Actor vs. Critic Networks
Figure 5: Critical depth and residual connections. Incrementally increasing depth results in marginal performance gains (left). However, once a critical threshold is reached, performance improves dramatically (right) for networks with residual connections.
Figure 6: Actor vs. Critic. In Arm Push Easy, scaling the critic is more effective; in Ant Big Maze, the actor matters more. For Humanoid, scaling both is necessary. These results suggest that actor and critic scaling can complement each other for CRL.
To investigate the role of scaling in the actor and critic networks, Figure˜6 presents the final performance for various combinations of actor and critic depths across three environments. Prior work (Nauman et al., 2024b, ; Lee et al., , 2024) focuses on scaling the critic network, finding that scaling the actor degrades performance. In contrast, while we do find that scaling the critic is more impactful in two of the three environments (Humanoid, Arm Push Easy), our method benefits from scaling the actor network jointly, with one environment (Ant Big Maze) demonstrating actor scaling to be more impactful. Thus, our method suggests that scaling both the actor and critic networks can play a complementary role in enhancing performance.
Deep Networks Unlock Batch Size Scaling
Figure 7: Deeper networks unlock batch size scaling. We find that as depth increases from 4 to 64 in Humanoid, larger networks can effectively leverage batch size scaling to achieve further improvements.
Scaling batch size has been well-established in other areas of machine learning (Chen et al., , 2022; Zhang et al., , 2024). However, this approach has not translated as effectively to reinforcement learning (RL), and prior work has even reported negative impacts on value-based RL (Obando-Ceron et al., , 2023). Indeed, in our experiments, simply increasing the batch size for the original CRL networks yields only marginal differences in performance (Figure˜7, top left).
To test this hypothesis, we study the effect of increasing the batch size for networks of varying depths. As shown in Figure˜7, scaling the batch size becomes effective as network depth grows. This finding offers evidence that by scaling network capacity, we may simultaneously unlock the benefits of larger batch size, potentially making it an important component in the broader pursuit of scaling self-supervised RL.
Training Contrastive RL with 1000+ Layers
We next study whether further increasing depth beyond 64 layers further improves performance. We use the Humanoid maze tasks as these are both the most challenging environments in the benchmark and also seem to benefit from the deepest scaling. The results, shown in Figure˜12, indicate that performance continues to substantially improve as network depth reaches 256 and 1024 layers in the Humanoid U-Maze environment. While we were unable to scale beyond 1024 layers due to computational constraints, we expect to see continued improvements with even greater depths, especially on the most challenging tasks.
Figure 8: We disentangle the effects of exploration and expressivity on depth scaling by training three networks in parallel: a “collector,” plus one deep and one shallow learner that train only from the collector’s shared replay buffer. In all three environments, when using a deep collector (i.e. good data coverage), the deep learner outperforms the shallow learner, indicating that expressivity is crucial when controlling for good exploration. With a shallow collector (poor exploration), even the deep learner cannot overcome the limitations of insufficient data coverage. As such, the benefits of depth scaling arise from a combination of improved exploration and increased expressivity working jointly.
4.5 Why Scaling Happens
Depth Enhances Contrastive Representations
@@ 4.5 Why Scaling Happens
The long-horizon setting has been a long-standing challenge in RL particularly in unsupervised goal-conditioned settings where there is no auxiliary reward feedback (Gupta et al., , 2019). The family of U-Maze environments requires a global understanding of the maze layout for effective navigation. We consider a variant of the Ant U-Maze environment, the U4-maze, in which the agent must initially move in the direction opposite the goal to loop around and ultimately reach it. As shown in Figure˜9, we observe a qualitative difference in the behavior of the shallow network (depth 4) compared to the deep network (depth 64). The visualized Q-values computed from the critic encoder representations reveal that the depth 4 network seemingly relies on Euclidean distance to the goal as a proxy for the Q value, even when a wall obstructs the direct path. In contrast, the depth 64 critic network learns richer representations, enabling it to effectively capture the topology of the maze as visualized by the trail of high Q values along the inner edge. These findings suggest that increasing network depth leads to richer learned representations, enabling deeper networks to better capture environment topology and achieve more comprehensive state-space coverage in a self-supervised manner.
Depth Enhances Exploration and Expressivity in a Synergized Way
Our earlier results suggested that deeper networks achieve greater state-action coverage. To better understand why scaling works, we sought to determine to whether improved data alone explains the benefits of scaling, or whether it acts in conjunction with other factors. Thus, we designed an experiment in Figure˜8 in which we train three networks in parallel: one network, the “collector," interacts with the environment and writes all experience to a shared replay buffer. Alongside it, two additional "learners", one deep and one shallow, train concurrently. Crucially, these two learners never collect their own data; they train only from the collector’s buffer. This design holds the data distribution constant while varying the model’s capacity, so any performance gap between the deep and shallow learners must come from expressivity rather than exploration. When the collector is deep (e.g., depth 32), across all three environments the deep learner substantially outperforms the shallow one across all three environments, indicating that the expressivity of the deep networks is critical. On the other hand, we repeat the experiment with shallow collectors (e.g., depth 4), which explores less effectively and therefore populates the buffer with low-coverage experience. Here, both the deep and shallow learners struggle and achieve similarly poor performance, which indicates that the deep network’s additional capacity does not overcome the limitations of insufficient data coverage. As such, scaling depth enhances exploration and expressivity in a synergized way: stronger learning capacity drives more extensive exploration, and strong data coverage is essential to fully realize the power of stronger learning capacity. Both aspects jointly contribute to improved performance.
Figure 10: We visualize state-action embeddings from shallow (depth 4) and deep (depth 64) networks along a successful trajectory in the Humanoid task. Near the goal, embeddings from the deep network expand across a curved surface, while those from the shallow network form a tight cluster. This suggests that deeper networks may devote greater representational capacity to regions of the state space that are more frequently visited and play a more critical role in successful task completion.
Deep Networks Learn to Allocate Greater Representational Capacity to States Near the Goal
In Figure˜10 we take a successful trajectory in the Humanoid environment and visualize the embeddings of state-action encoder along this trajectory for both deep vs. shallow networks. While the shallow network (Depth 4) tends to cluster near-goal states tightly together, the deep network produces more "spread out" representations. This distinction is important: in a self-supervised setting, we want our representations to separate states that matter—particularly future or goal-relevant states—from random ones. As such, we want to allocate more representational capacity to such critical regions. This suggests that deep networks may learn to allocate representational capacity more effectively to state regions that matter most for the downstream task.
Deeper Networks Enable Partial Experience Stitching
@@
Another key challenge in reinforcement learning is learning policies that can generalize to tasks unseen during training. To evaluate this setting, we designed a modified version of the Ant U-Maze environment. As shown in Figure˜11 (top right), the original JaxGCRL benchmark assesses the agent’s performance on the three farthest goal positions located on the opposite side of the wall. However, instead of training on all possible subgoals (a superset of the evaluation state-goal pairs), we modified the setup to train on start-goal pairs that are at most 3 units apart, ensuring that none of the evaluation pairs ever appear in the training set. Figure˜11 demonstrates that depth 4 networks show limited generalization, solving only the easiest goal (4 units away from the start). Depth 16 networks achieve moderate success, while depth 64 networks excel, sometimes solving the most challenging goal position. These results suggest that the increasing network depth results in some degree of stitching, combining $\leq$3-unit pairs to navigate the 6-unit span of the U-Maze.
The (CRL) Algorithm is Key
In Appendix A, we show that scaled CRL outperforms other baseline goal-conditioned algorithms and advance the SOTA for goal-conditioned RL. We observe that for temporal difference methods (SAC, SAC+HER, TD3+HER), the performance saturates for networks of depth 4, and there is either zero or negative performance gains from deeper networks. This is in line with previous research showing that these methods benefit mainly from width (Lee et al., , 2024; Nauman et al., 2024b, ). These results suggest that the self-supervised CRL algorithm is critical.
We also experiment with scaling more self-supervised algorithms, namely Goal-Conditioned Behavioral Cloning (GCBC) and Goal-Conditioned Supervised Learning (GCSL). While these methods yield zero success in certain environments, they show some utility in arm manipulation tasks. Interestingly, even a very simple self-supervised algorithm like GCBC benefits from increased depth. This points to a promising direction for future work of further investigating other self-supervised methods to uncover potentially different or complementary recipes for scaling self-supervised RL.
Finally, recent work has augmented goal-conditioned RL with quasimetric architectures, leveraging the fact that temporal distances satisfy a triangle inequality–based invariance. In Appendix A, we also investigate whether the depth scaling effect persists when applied to these quasimetric networks.